# Extracting Phonon Quasiparticles from Molecular Dynamics Simulations

### v2

@@ Abstract
Phonon anharmonicity is ubiquitous in real materials and is crucial for understanding thermal properties and phase stability. In this work, we show that phonon quasiparticles are optimally described by modes with maximum lifetimes, and prove that information about these quasiparticles is contained in two small matrices $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from molecular dynamics simulations. Based on these knowledge, we proposed an optimization scheme, which allows us to efficiently determine temperature-dependent phonon modes along with their frequencies and lifetimes. We verified this method by applying it to silicon and cubic CaSiO3, where it successfully captured their temperature-dependent phonon behaviors and the well-known phonon softening in cubic CaSiO3. This theory provides a convenient tool for investigating phonon quasiparticles and can be extended to study other quasiparticles, such as electrons, holes, and magnons.
@@
Over the past decades, researchers have developed various methods to study phonons beyond the harmonic approximation. A natural method is to include phonon-phonon collisions using the third-order or even fourth-order force constants [9, 10, 11, 12, 13, 14, 15, 16]. This approach has effectively determined phonon lifetimes across various materials, including semiconductors [17, 18, 19, 20], rocksalt materials[21], zincblende compounds[22, 21, 23, 24] and wurtzite compounds [24], and two-dimensional crystals [25, 26, 27, 28]. Although these perturbative approaches have proven successful, numerous non-perturbative methods have also emerged, such as the self-consistent harmonic approximation (SCHA)[29, 30, 31, 32], self-consistent ab initio lattice dynamics (SCAILD) [33], stochastic self-consistent harmonic approximation (SSCHA)[34], self-consistent phonon calculations (SCPH) [35], Wigner transport equation (WTE)[36, 37] and quantum self-consistent ab initio lattice dynamics (QSCHILD) [38]. Additionally, molecular dynamics (MD), especially the ab initio molecular dynamics (AIMD), offers another powerful non-perturbative approach for studying phonon anharmonicity, as atomic trajectories naturally incorporate anharmonic effects to all orders. The temperature-dependent effective potential (TDEP) method[39] extracts effective force constants from AIMD simulations, enabling the calculation of anharmonic phonon properties. Similarly, the phonon quasiparticle approach (PQA) projects AIMD trajectories onto normal modes directly obtain anharmonic phonon frequencies and lifetimes[40, 41, 42]. With all these methods yielding phonon quasiparticles via distinct approaches, a fundamental question arises: How accurate are these phonon quasiparticles in modeling the vibration of a real system?
@@
The key is lifetime. In fact, phonon quasiparticles with inaccurate phonon modes typically exhibit shorter lifetimes. Therefore, modes with maximized lifetimes provide the most accurate representation of of phonon quasiparticles. This principle inspires a concise and efficient method, developed in this study, to extract phonon quasiparticles from MD simulations.
@@
For a pure harmonic system with $n$ atoms in primitive cell, the $j$-th phonon mode at $\vec{k}$ point of Brillouin zone is $\varphi_{i\sigma}^{\vec{k},j}$ ($i=1,2,..,n;\,\sigma=x,y,z$). There are 3nn phonon modes in total, forming a complete orthonormal basis set. Then, the atom trajectories or velocities in a MD simulation can be expanded as $\mathcal{R}_{i\sigma}^{I}(t)=Re\left(\sum_{\vec{k},j}c^{\vec{k},j}\varphi_{i\sigma}^{\vec{k},j}e^{\mathbf{i}(\vec{k}\cdot\vec{R}_{I}-\omega_{\vec{k},j}t)}\right)$, where $\omega$ is eigen frequency, $t$ is time, and $\vec{R}_{I}$ is the displacement vector of the $I$-th cell. The coefficient $c_{\vec{k},j}$ included here represents the weight of mode $\varphi^{\vec{k},j}$. Since $\mathcal{R}_{i\sigma}^{I}(t)$ contains phonon modes at all $\vec{k}$ points, we can filter out the modes at specific $\vec{k}$ by real space Fourier transformation:
@@
Eq. 2 also applies to real systems. The key differences are that both modes and frequencies depend on MD temperature, and the coefficients $c^{\vec{k},j}$ become time-dependent due to mode scattering caused by anharmonicity. A non-zero $|\frac{dc^{\vec{k},j}(t)}{dt}|$ leads to a finite quasiparticle lifetime, which further depends on the choose of modes. If inaccurate modes and the associated frequencies are used to expand MD trajectories or velocities, the coefficients $c^{\vec{k},j}(t)$ will vary more rapidly over time. This suggests that optimal phonon quasiparticles can be defined as those modes that minimize $|\frac{dc^{\vec{k},j}(t)}{dt}|$, equivalently speaking maximize quasiparticle lifetimes.
Next, we describe the method for obtaining optimal phonon quasiparticles. After a MD simulation, the column vector $g^{\vec{k}}(t)$ is constructed directly from atom trajectories, velocities, or their mass-weighted forms. The time-dependent coefficient is then given by $c^{\vec{k},j}(t)=2(\varphi^{\vec{k},j})^{+}g^{\vec{k}}(t)e^{\mathbf{i}\omega_{\vec{k},j}t}$. For mathematical convenience, we minimize $|\frac{dc^{j}(t)}{dt}|^{2}$ rather than $|\frac{dc^{\vec{k},j}(t)}{dt}|$ to optimize phonon quasiparticles. Since the optimization was performed individually for each $\vec{k}$ point, for brevity, we omit the $\vec{k}$ superscript in the following discussions, unless needed for clarity.
@@
where $t_{i}$ and $t_{f}$ are the initial and final simulation times, and $\mathcal{G}^{j}=\mathcal{G}_{0}+\omega_{j}\mathcal{P}+\omega_{j}^{2}\mathcal{Q}$. The matrices $\mathcal{G}_{0}$, $\mathcal{P}$, and $\mathcal{Q}$ here are independent of the initial guessed phonon modes. They are all Hermitian and can be constructed solely from the MD trajectory,
@@
The central equations of this method, Eqs. 3 and 4, rely on the matrices $\mathcal{S}$ and $\mathcal{Q}$, which are constructed directly from molecular MD trajectories or velocities. The procedure begins by diagonalizing $\mathcal{Q}$ to obtain its eigenvectors, which serve as the initial phonon modes. These modes are then optimized by constructing the matrices $S$ and $Q$ and applying Eqs. 3 and 4, yielding the optimal phonon modes and frequencies. Finally, the linewidths are calculated from ACF.
@@
$W(\omega)=\frac{1}{\pi}\sum_{j}\frac{1}{e^{\hbar\omega_{j}/kT}-1}\frac{\Gamma_{j}}{\hbar(\omega-\omega_{j})^{2}+\Gamma^{2}},$
@@
Figure 3: Phonon dispersions for CaSiO3 at (a)400 K and (b) 1000 K. The linewidths are represented by the thickness of the curves. Panels (c) and (d) display the corresponding phonon spectral functions.
@@
Unlike silicon, cubic CaSiO3 is a typical polar material. In such systems, the transverse optical (TO) phonon modes and longitudinal optical (LO) phonon modes split due to long-range Coulomb interactions. Therefore, after obtaining the effective force constants, we need to apply corrections[57, 58] based on Born effective charges and the dielectric tensor. The lattice parameter of the unit cell was set to 3.45 Å, corresponding to the pressure at the top of the lower mantle. The calculated phonon dispersions of cubic CaSiO3 at 400 K and 1000 K are shown in Fig. 3(a,b). The LO and TO phonons split at the $\Gamma$ point. The lowest three branches of phonons around the $R$ point have zero frequencies, corresponding to the imaginary frequencies in previous studies—indicating that cubic CaSiO3 is unstable at 400 K. However, these phonons have positive frequencies at 1000 K, showing that the material becomes stable at higher temperatures. The top three phonon branches throughout the entire Brillouin zone and the lowest three phonon branches at the $R$ point exhibit relatively large linewidths. The corresponding phonon spectral functions are shown in Fig. 3(c,d).
The effectiveness of this method stems from the fact that phonon quasiparticles govern atomic motion in crystals, with key information stored in compact matrices. This paradigm is broadly applicable to other systems governed by their own quasiparticles. For example, in the case of magnons, compact matrices can be derived from spin trajectories calculated using the Landau-Lifshitz-Gilbert equation. Similarly, for electrons and holes, compact matrices or kernel functions can be extracted from time-dependent DFT calculation results. The corresponding quasiparticles can then be systematically determined from these matrices or kernel functions.
In summary, this work defines optimal phonon quasiparticles as those that maximize their lifetimes. We demonstrate that all information about these quasiparticles is stored in two simple matrices, $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from MD trajectories or velocities. The eigen vectors of $\mathcal{Q}$ can be used as the initial phonon modes. Then the optimal phonons and their corresponding frequencies and lifetimes can be obtained from an optimization scheme. Tests on silicon show weak anharmonicity with temperature-dependent shifts, while cubic CaSiO3 exhibits strong anharmonic effects, being unstable at 400 K but stabilized at 1000 K. These tests indicate the reliability of this method. This theory and optimization method provide a clear picture of phonon quasiparticles in real materials, offering a convenient tool for their investigation. Its core principles are readily transferable, paving the way for the study of a broad spectrum of other quasiparticles.
@@ Acknowledgements.
W. Li and Y. Lu contributed equally to this work. This work was supported by the National Natural Science Foundation of China under Grants 12374054 and 12022415.

### v4

# Correlation Matrix Method for Phonon Quasiparticles

@@ Abstract
Phonon anharmonicity is ubiquitous in real materials and is crucial for understanding thermal properties and phase stability. In this work, we show that anharmonic phonon modes can be obtained by maximizing their vibration stability during fitting the atomic trajectory. We prove that all information about these quasiparticles is contained in two small correlation matrices $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from molecular dynamics simulations. Based on these matrices, we proposed an optimization scheme, which allows us to efficiently determine temperature-dependent phonon modes along with their frequencies and lifetimes. We verified this method by applying it to silicon and cubic CaSiO3, where it successfully captured their temperature-dependent phonon behaviors and the well-known phonon softening in cubic CaSiO3. This theory provides a convenient tool for investigating phonon quasiparticles and can be extended to study other quasiparticles, such as electrons, holes, and magnons.
@@
Over the past decades, researchers have developed various methods to study phonons beyond the harmonic approximation. A natural method is to include phonon-phonon collisions using the third-order or even fourth-order force constants [Deinzer, 12, 15]. This approach has effectively determined phonon lifetimes and thermal conductivity across various materials, including semiconductors [18, 19, 20], rocksalt materials[21], zincblende compounds[22, 21, 23, 24] and wurtzite compounds [24], and two-dimensional crystals [25, 26, 27, 28]. Although these perturbative approaches have proven successful, numerous non-perturbative methods have also emerged, such as the self-consistent harmonic approximation (SCHA)[29, 30, 31, 32], self-consistent ab initio lattice dynamics (SCAILD) [33], stochastic self-consistent harmonic approximation (SSCHA)[34], self-consistent phonon calculations (SCPH) [35], quantum self-consistent ab initio lattice dynamics (QSCHILD) [38], and use compressive sensing lattice dynamics[16]. Moreover, classical and ab initio molecular dynamics (MD) offers another powerful non-perturbative approach for studying phonon anharmonicity, as atomic trajectories naturally incorporate anharmonic effects to all orders. The fluctuation-dissipation theorem can be used to obtain dynamical matrices from MD simulations[39]. The temperature-dependent effective potential (TDEP) method[40, 41] and temperature dependent force constants[42, 43, 9] can be extracted from AIMD simulations, enabling the calculation of anharmonic phonon properties. Similarly, the phonon quasiparticle approach (PQA) projects AIMD trajectories onto normal modes to directly obtain anharmonic phonon frequencies and lifetimes[44, 45, 46]. With all these methods yielding phonon quasiparticles via distinct approaches, a fundamental question arises: How accurate are these phonon quasiparticles in modeling the vibration of a real system?
@@
The key is vibration stability. In fact, phonon quasiparticles with inaccurate phonon modes exhibit less vibration stability when decomposing atom trajectories. Therefore, modes with maximized stability provide the most accurate representation of phonon quasiparticles. This principle inspires a concise and efficient method, developed in this study, to extract phonon quasiparticles from MD simulations.
@@
For a pure harmonic system with $n$ atoms in primitive cell, the $j$-th phonon mode at $\vec{k}$ point of Brillouin zone is $\varphi_{i\sigma}^{\vec{k},j}$ ($i=1,2,..,n;\,\sigma=x,y,z$). There are 3nn phonon modes in total, forming a complete orthonormal basis set. Then, the mass-weighted atom trajectories or velocities in a MD simulation can be expanded as $\mathcal{R}_{i\sigma}^{I}(t)=Re\left(\sum_{\vec{k},j}c^{\vec{k},j}\varphi_{i\sigma}^{\vec{k},j}e^{\mathbf{i}(\vec{k}\cdot\vec{R}_{I}-\omega_{\vec{k},j}t)}\right)$, where $\omega$ is eigen frequency, $t$ is time, and $\vec{R}_{I}$ is the displacement vector of the $I$-th cell. The coefficient $c_{\vec{k},j}$ represents the weight of mode $\varphi^{\vec{k},j}$. Since $\mathcal{R}_{i\sigma}^{I}(t)$ contains phonon modes at all $\vec{k}$ points, we can filter out the modes at specific $\vec{k}$ by real space Fourier transformation:
@@
Eq. 2 also applies to real systems. The key differences are that both modes and frequencies depend on MD temperature, and the coefficients $c^{\vec{k},j}$ become time-dependent due to mode scattering caused by anharmonicity. The value $|\frac{dc^{\vec{k},j}(t)}{dt}|$ quantifies the vibration stability of mode $\varphi^{\vec{k},j}$. If inaccurate modes and the associated frequencies are used to expand MD trajectories, the coefficients $c^{\vec{k},j}(t)$ will vary more rapidly over time. This is clearer in a pure harmonic system, when MD trajectories are decomposed by non-eigenvectors, the normally invariant coefficients become time-dependent. This suggests that optimal phonon quasiparticles can be defined as those modes that minimize $|\frac{dc^{\vec{k},j}(t)}{dt}|$, equivalently speaking maximize stability.
Next, we describe the method for obtaining optimal phonon quasiparticles. After a MD simulation, the column vector $g^{\vec{k}}(t)$ is constructed directly from either mass-weighted atom trajectories or mass-weighted velocities. The time-dependent coefficient is then given by $c^{\vec{k},j}(t)=2(\varphi^{\vec{k},j})^{+}g^{\vec{k}}(t)e^{\mathbf{i}\omega_{\vec{k},j}t}$. For mathematical convenience, we minimize $|\frac{dc^{j}(t)}{dt}|^{2}$ rather than $|\frac{dc^{\vec{k},j}(t)}{dt}|$ to optimize phonon quasiparticles. Since the optimization was performed individually for each $\vec{k}$ point, for brevity, we omit the $\vec{k}$ superscript in the following discussions, unless needed for clarity.
@@
where $t_{i}$ and $t_{f}$ are the initial and final simulation times, and $\mathcal{G}^{j}=\mathcal{G}_{0}+\omega_{j}\mathcal{P}+\omega_{j}^{2}\mathcal{Q}$. The matrices $\mathcal{G}_{0}$, $\mathcal{P}$, and $\mathcal{Q}$ here are independent of the initial guessed phonon modes. They are Hermitian and can be constructed solely from the MD trajectory as equal-time correlation matrices,
@@
The phonon quasiparticle calculation procedure begins with constructing $\mathcal{S}$ and $\mathcal{Q}$ from MD trajectories. Next, $\mathcal{Q}$ is diagonalized to obtain its eigenvectors, which serve as the initial phonon modes. These modes are then optimized by applying Eqs. 3 and 4, yielding the optimal phonon modes and frequencies. Finally, the linewidths are calculated from ACF.
@@
$W(\omega)=\frac{1}{\pi}\sum_{j}\frac{1}{e^{\hbar\omega_{j}/kT}-1}\frac{\Gamma_{j}}{(\omega-\omega_{j})^{2}+\Gamma^{2}},$
@@
Figure 3: Phonon dispersions and modes for CaSiO3 at (a)400 K and (b) 1000 K. The linewidths are represented by the line thickness. Panels (d, e) display the corresponding phonon spectral functions. Panels (c, f) show the lowest phonon modes ML and RL at 1000 K.
@@
Unlike silicon, cubic CaSiO3 is a typical polar material. In such systems, the transverse optical (TO) phonon modes and longitudinal optical (LO) phonon modes split due to long-range Coulomb interactions. Therefore, after obtaining the effective force constants, we need to apply corrections[61, 62] based on Born effective charges and the dielectric tensor. The lattice parameter of the unit cell was set to 3.45 Å, corresponding to the pressure at the top of the lower mantle. The calculated phonon dispersions of cubic CaSiO3 at 400 K and 1000 K are shown in Fig. 3(a,b). The resulting phonon spectral functions are shown in Fig. 3(d,e). The LO and TO phonons split at the $\Gamma$ point. The lowest three branches of phonons around the $R$ point have zero frequencies, corresponding to the imaginary frequencies in previous studies—indicating that cubic CaSiO3 is unstable at 400 K. But their frequencies are positive at 1000 K, showing that the material becomes stable at higher temperatures. The lowest phonon mode at $M$ point also shows soften at low temperature. The lowest modes at the $M$ and $R$ points, shown in 3(c,f), are characterized by the vibration of O atoms.
The use of compact matrices ensures the efficiency of this method, while the absence of approximations guarantees its accuracy. This paradigm is broadly applicable to other systems governed by their own quasiparticles. For example, in the case of magnons, compact matrices can be derived from spin trajectories calculated using the Landau-Lifshitz-Gilbert equation. Similarly, for electrons and holes, compact matrices or kernel functions can be extracted from time-dependent DFT results. The corresponding quasiparticles can then be systematically determined from these matrices or kernel functions.
In summary, this work defines optimal phonon quasiparticles as those that maximize their stability. We demonstrate that all information about these quasiparticles is stored in two simple matrices, $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from MD trajectories or velocities. The eigen vectors of $\mathcal{Q}$ can be used as the initial phonon modes. Then the optimal phonons and their corresponding frequencies and lifetimes can be obtained from an optimization scheme. Tests on silicon show weak anharmonicity with temperature-dependent shifts, while cubic CaSiO3 exhibits strong anharmonic effects, being unstable at 400 K but stabilized at 1000 K. These tests show the reliability of this method. This theory and optimization method provide a clear picture of phonon quasiparticles in real materials, offering a convenient tool for their investigation. Its core principles are readily transferable, paving the way for the study of a broad spectrum of other quasiparticles.
@@ Acknowledgements.
W. Li and Y. Lu contributed equally to this work. We thank T. Sun, F. Zhou, Z.-K. Liu, A. Togo, and S. S. Naghavi for useful discussions. This work was supported by the National Natural Science Foundation of China under Grants 12374054 and 12022415. We acknowledge the computing resources of the Tencent TEFS platform (https://tefscloud.com).